# 2Preon Health Oy, Finland

### v1

@@ Abstract
One of the primary challenges in brain tumor segmentation arises from the uncertainty of voxels close to tumor boundaries. However, the conventional process of generating ground truth segmentation masks fails to treat such uncertainties properly. Those “hard labels” with 0s and 1s conceptually influenced the majority of prior studies on brain image segmentation. As a result, tumor segmentation is often solved through voxel classification. In this work, we instead view this problem as a voxel-level regression, where the ground truth represents a certainty mapping from any pixel based on the distance to tumor border. We propose a novel ground truth label transformation, which is based on a signed geodesic transform, to capture the uncertainty in brain tumors’ vicinity, while maintaining a margin between positive and negative samples. We combine this idea with a Focal-like regression L1-loss that enables effective regression learning in high-dimensional output space by appropriately weighting voxels according to their difficulty. We thoroughly conduct an experimental evaluation to validate the components of our proposed method, compare it to a diverse array of state-of-the-art segmentation models, and show that it is architecture-agnostic. The code of our method is made publicly available at https://github.com/Oulu-IMEDS/SiNGR/.
@@ 1 Introduction
In the existing literature on brain tumor segmentation, the first line of research aims to improve the effectiveness of DL architectures by increasing the model capacity and/or embedding domain knowledge into architecture design [7, 16, 26, 23, 8]. The second line of research studies novel segmentation losses that are preferably correlated with metrics of interest such as Dice score or Intersection-of-Union (IoU) [3, 2, 15, 25, 24]. Finally, the third group of studies investigates alternative ways to define the ground truth (GT) masks through soft labels [20, 12]. Our study is at the intersection of the last two directions, as we aim to define the GT masks through soft labels, as well as to develop a new loss function for the problem.
@@
We observe that the complexity of the segmentation annotation process is in the labeling of voxels around the boundaries of objects of interest (OOI). The uncertainty of a voxel’s label is proportional to the inverse of its distance to the nearest boundary, as well as the visual blurriness in this region. Although unsigned GeoDT naturally allows us to capture these properties, additional signals are needed to differentiate foreground from background voxels. Moreover, voxels significantly distant from the OOI exhibit very low uncertainty; thus, these voxels should be marked in a manner that directs the model to pay less attention to them.
@@ 2.2 Signed Normalized Geodesic Transform
For signed GeoDT, one typically runs the GeoDT transform twice for the foreground and background regions, which is intensively costly [6]. To speed up the signed GeoDT, we merely consider a set $R_{B}$ of boundary voxels of OOI extracted from a 0-1 segmentation mask $M$ using the Canny edge detector [5]. Hereinafter, as $R_{B}$ is our primary region of interest, we omit $R_{B}$ and $I$ from Eq. 1 for simplicity, that is $G^{\lambda}_{i}=G^{\lambda}(i;R_{B},I)$. Given $R_{B}$, we apply unsigned GeoDT for the whole image to produce an unsigned map. Afterwards, we rely on $M$ to specify the signs of the obtained map, that is $s_{i}G^{\lambda}_{i},\forall i\in\Omega$, where $s_{i}=\mathrm{sign}(2M_{i}-1)$ is the sign of voxel $i$. As the uncertainty of human annotations is primarily around the boundary $R_{B}$, we thus ignore regions substantially far from the boundary. As such, we let $\lambda=0$, and define the neighboring region of the boundary
@@
where $\tau=\max_{j\in\mathcal{B}}G^{\lambda}_{j}$ is a normalization constant for each pair $(I,M)$, and $\delta\in[0,1)$ is a margin hyperparameter. The margin $\delta$ is needed to ensure the discrepancy in the signed normalized map between foreground and background voxels.
@@ 2.3 Focal-L1 Loss
where $\varepsilon$ is a positive constant to avoid numerical issues, $\gamma$ is a positive hyperparameter, and $\mathbb{I}(\cdot)$ is the indicator function. A graphical comparison between our loss and the L1 loss is presented in Figure 2. Note that the backpropagation is not applied to the weighting term. Whereas the weighting term’s numerator is fixed at $1$ for hard cases ($\mathcal{S}_{i}\mathcal{Z}_{i}<0$), it is scaled down to less than $1$ for easy ones ($\mathcal{S}_{i}\mathcal{Z}_{i}\geq 0$). Meanwhile, the denominator reduces the importance of highly certain voxels (with high absolute value), as they are straightforward to predict. Overall, the sample weighting helps the loss to prioritize enforcing difficult voxels over simple ones.
@@ 3.1 Setup
Implementation details. All models were trained on Nvidia V100 GPUs. We implemented our pipeline using Pytorch and followed a standard data processing configuration for all methods. We used the FastGeodist library to compute unsigned GeoDT maps [1]. During the training process on BraTS, the 3D MR images were randomly cropped to $128\times 128\times 128$ cubes, and augmented by flipping, intensity scaling, and intensity shifting. We performed the window-slicing technique with a window size of $128\times 128\times 128$ in the validation and test stages. We utilized the Adam optimizer with an initial learning rate of $10^{-4}$, a weight decay of $10^{-5}$, and a batch size of $2$. Regarding SiNGR-specific hyperparameters, we empirically found $\lambda=0.5$ and $\gamma=1$ working the best. The hyperparameter $\varepsilon$ in Eq. 6 was set to $0.0$. We performed 00-thresholding to binarize our predicted maps. For performance evaluation, we utilized image-wise IoU, Dice score, and $95\%$ Hausdorff distance (HD95).
We compared our method to a diverse array of references specializing in 3D data as listed in Tables 1 and 2. In addition, we validated the SiNG transform and the Focal-L1 loss against other soft-label baselines – namely label smoothing (LS) [18] and Geodesic label smoothing (GeoLS) [20] – together with Jaccard metric loss (JML) [24], which specializes in optimizing with soft labels.
@@ 3.2 Results
LGG FLAIR dataset. We present the detailed results on the LGG FLAIR test set in Table 2. On this small dataset, the impact of SiNGR was particularly significant on UNet3D, and this setting also acquired the highest IoU and Dice score. Compared to the UNet3D reference, our method resulted in significant gains of $13.5\%$ IoU, $13.2\%$ Dice score, and $22.2mm$ HD95. Using our method also consistently led to substantial improvements for Swin-UNETR and TransBTS (see Figures 3c and 3d). The qualitative results are presented in Suppl. Figure S2.
Impact of SiNG and Focal-L1. We investigated the effects of the components of our method and demonstrated the results on the BraTS test set in Table 3. Two unsigned label smoothing techniques [18, 20] did not perform well on the task. Compared to LS, our method outperformed with differences of $8.4\%$ and $7.8mm$ in average Dice score and average HD95, respectively. Additionally, among L1, L2, and product [27] losses, L1 was the best combination with the SiNG transform. However, L1 setting achieved an average Dice of $1.1\%$ lower than our method. Moreover, the empirical evidence showed the importance of the margin in the SiNG transform as well as the sample weighting coefficient in Focal-L1 loss. Notably, excluding the margin $\delta$ in SiNG led to a substantial drop of $3.4\%$ average Dice score and an increase of $1.5mm$ average HD95.
@@
4 Conclusions
We have introduced a simple approach to segmentation of brain tumors through voxel-wise regression.We proposed the novel SiNG transform that allows us to convert 0-1 annotated masks to soft labels that take into account the uncertainty of the labeling process. In addition, we have introduced the Focal-L1 loss to effectively weight voxels according to their difficulty. Our empirical findings indicate that our method consistently enhances performance across different DL architectures. We will make our implementation publicly accessible to the community.

### v4

# SiNGR: Brain Tumor Segmentation via Signed Normalized Geodesic Transform Regression

@@ Abstract
One of the primary challenges in brain tumor segmentation arises from the uncertainty of voxels close to tumor boundaries. However, the conventional process of generating ground truth segmentation masks fails to treat such uncertainties properly. Those “hard labels” with 0s and 1s conceptually influenced the majority of prior studies on brain image segmentation. As a result, tumor segmentation is often solved through voxel classification. In this work, we instead view this problem as a voxel-level regression, where the ground truth represents a certainty mapping from any pixel to the border of the tumor. We propose a novel ground truth label transformation, which is based on a signed geodesic transform, capturing the uncertainty of brain tumors’ vicinity. We combine this idea with a Focal-like regression L1-loss that enables effective regression learning in high-dimensional output space by appropriately weighting voxels according to their difficulty. We thoroughly conduct an experimental evaluation, validating the components of our proposed method, comparing it to a diverse array of state-of-the-art segmentation models, and showing that it is architecture-agnostic. The code of our method is made publicly available (https://github.com/Oulu-IMEDS/SiNGR/).
@@ 1 Introduction
In the existing literature on brain tumor segmentation, the first line of research aims to improve the effectiveness of DL architectures by increasing the model capacity and/or embedding domain knowledge into architecture design [7, 17, 27, 24, 8]. The second line of research studies novel segmentation losses that are preferably correlated with metrics of interest such as Dice score or Intersection-of-Union (IoU) [3, 2, 16, 26, 25]. Finally, the third group of studies investigates alternative ways to define the ground truth (GT) masks through soft labels [21, 12]. Our study is at the intersection of the last two directions, as we aim to define the GT masks through soft labels, as well as to develop a new loss for the problem.
@@
We observe that the complexity of the segmentation annotation process is in the labeling of voxels around the boundaries of objects of interest (OOI). The uncertainty of a voxel’s label is proportional to its distance to the nearest boundary, as well as the visual blurriness in this region. Although unsigned GeoDT naturally allows us to capture these properties, additional signals are needed to differentiate foreground from background voxels. Moreover, voxels significantly distant from the OOI exhibit very low uncertainty; thus, these voxels should be marked in a manner that directs the model to pay less attention to them.
@@ 2.2 Signed Normalized Geodesic Transform
For signed GeoDT, one typically runs the GeoDT transform twice for the foreground and background regions, which is intensively costly [6]. To speed up the signed GeoDT, we merely consider a set $R_{B}$ of boundary voxels of OOI extracted from a 0-1 segmentation mask $M$ using the Canny edge detector [5]. Hereinafter, as $R_{B}$ is our primary region of interest, we omit $R_{B}$ and $I$ from Eq. 1 for simplicity, that is $G^{\lambda}_{i}=G^{\lambda}(i;R_{B},I)$. Given $R_{B}$, we apply unsigned GeoDT for the whole image to produce an unsigned map. Afterwards, we rely on $M$ to specify the signs of the obtained map, that is $s_{i}G^{\lambda}_{i},\forall i\in\Omega$, where $s_{i}=\mathrm{sign}(2M_{i}-1)$ is the sign of voxel $i$.
As the uncertainty of human annotations primarily is around the boundary $R_{B}$, we thus ignore regions substantially far from the boundary. As such, we let $\lambda=0$, and define the neighboring region of the boundary
@@
where $\tau=\max_{j\in\mathcal{B}}G^{\lambda}_{j}$ is a normalization constant for each pair $(I,M)$, and $\delta\in[0,1)$ is a margin hyperparameter. The margin $\delta$ is needed to ensure the discrepancy in the signed normalized map between foreground and background pixels.
@@ 2.3 Focal-L1 Loss
where $\varepsilon$ is a positive constant to avoid numerical issues, $\gamma$ is a positive hyperparameter, and $\mathbb{I}(\cdot)$ is the indicator function. A graphical comparison between our loss and the L1 loss is presented in Figure 2. Note that the backpropagation is not applied to the weighting term. Whereas the weighting term’s numerator is fixed at $1$ for hard cases ($\mathcal{S}_{i}\mathcal{Z}_{i}<0$), it is scaled down to less than $1$ for easy ones ($\mathcal{S}_{i}\mathcal{Z}_{i}\geq 0$). Meanwhile, the denominator reduces the importance of highly certain voxels (with high absolute value), as they are straightforward to predict. Overall, the sample weighting helps the loss to prioritize enforcing hard voxels over simple ones.
@@ 3.1 Setup
Implementation details. All models were trained on Nvidia V100 GPUs. We implemented our pipeline using Pytorch and followed a standard data processing configuration for all methods. We used the FastGeodist library to compute unsigned GeoDT maps [1]. During the training process on BraTS, the 3D MR images were randomly cropped to $128\times 128\times 128$ cubes, and augmented by flipping, intensity scaling, and intensity shifting. We performed the window-slicing technique with a window size of $128\times 128\times 128$ in the validation and test stages. The employed patch size on LGG was $128\times 128\times 32$. We utilized the Adam optimizer with an initial learning rate of $10^{-4}$, a weight decay of $10^{-5}$, and a batch size of $2$. Regarding SiNGR-specific hyperparameters, we empirically found $\lambda=0.5$, $\delta=0.5$, and $\gamma=1$ working the best. The hyperparameter $\varepsilon$ in Eq. 6 was set to $0.0$. We performed 00-thresholding to binarize our predicted maps. For performance evaluation, we utilized image-wise IoU, Dice score, and $95\%$ Hausdorff distance (HD95).
We compared our method to a diverse array of references specializing in 3D data as listed in Tables 1 and 2. In addition, we validated the SiNG transform and the Focal-L1 loss against other soft-label baselines – namely label smoothing (LS) [19] and Geodesic label smoothing (GeoLS) [21] – together with the Jaccard Metric Loss (JML) [25], which specializes in optimizing with soft labels.
@@ 3.2 Results
LGG FLAIR dataset. We present the detailed results on the LGG FLAIR test set in Table 2. On this dataset, the impact of SiNGR was particularly significant on UNet3D, and this also acquired the highest IoU and Dice score. Compared to the UNet3D reference, our method resulted in significant gains of $13.5\%$ IoU, $13.2\%$ Dice score, and $22.2mm$ HD95. Using our method also consistently led to substantial improvements for Swin-UNETR and TransBTS (see Figures 3c and 3d). The qualitative results are presented in Suppl. Figure S2.
Impact of SiNG and Focal-L1. We investigated the effects of the components of our method and demonstrated the results on the BraTS test set in Table 3. Two unsigned label smoothing techniques [19, 21] did not perform well on the task. Compared to LS with JML, our method outperformed with differences of $8.4\%$ and $7.8mm$ in average Dice score and average HD95, respectively. Additionally, among L1, L2, and product [28] losses, L1 was the best combination with the SiNG transform. However, that setting achieved an average Dice of $1.1\%$ lower than our method. Moreover, the empirical evidence showed the importance of the margin in the SiNG transform as well as the sample weighting coefficient in Focal-L1 loss. Notably, excluding the margin $\delta$ led to a substantial drop of $3.4\%$ average Dice score and an increase of $1.5mm$ average HD95.
@@
4 Conclusion
We have introduced a simple approach to segmentation of brain tumors through voxel-wise regression. We proposed the novel SiNG transform that allows us to convert 0-1 annotated masks to soft labels that take into account the uncertainty of the labeling process. In addition, we have introduced the Focal-L1 loss to effectively weight voxels according to their difficulty. Our empirical findings indicate that our method consistently enhances performance across different DL architectures.